A pendulum that weighs 1.0 kg is pulled sideways so that it is raised a vertical distance of 3 m above its resting position. Find the maximum speed the pendulum reaches after being released.


1
Expert's answer
2021-02-10T10:06:30-0500

According to the conservation energy law, the maximum potential energy of the pendulum is equal to its maximum kinetic energy. Writing down the expressions for each energy, obtain:


"mgh_{max} = \dfrac{mv^2_{max}}{2}"

where "m" is the mass of the pendulum, "h_{max} = 3m" is the initial (and maximum) height, "g = 9.81m\/s^2" is the gravitational acceleration, and "v_{max}" is the maximum speed. Thus, obtain:


"v_{max} = \sqrt{2gh_{max}}\\
v_{max} = \sqrt{2\cdot 9.81\cdot 3} \approx 7.67m/s"

Answer. 7.67 m/s.